# Local wind speed forecasting at short time horizons relying on both Numerical Weather Prediction and observations from surrounding stations

### v1

@@ 3.4 Training, validation and test sets, model’s hyper-parameters
3.5 Metrics
In order to assess the forecasting performance of a model, the most commonly used metrics is the so-called Root-Mean-Squared-Error (RMSE), that is defined as:
$\text{RMSE}=\sqrt{\frac{1}{NH}\sum_{k=1}^{N}\sum_{i=1}^{H}\left({\widehat{V}}^{s_{k}}_{d_{k},t_{k}+h_{i}}-V^{s_{k}}_{d_{k},t_{k}+h_{i}}\right)^{2}}$ (11)
where $k$ belongs to either the validation or test subset. In any case, $N$ represents the number of samples of the considered subset and $H$ stands for the number of prediction horizons ($H=6$ in our study). RMSE corresponds to the empirical deviation (in $m\;s^{-1}$) of observed velocity as respect to the prediction averaged over all stations, all days, all times and all horizons. Let us emphasize that, since our goal is to compare models providing single-point predictions, we treat model ${\mathcal{M}}^{\prime}$ that provides a probabilistic forecasting as an intermediate step for generating deterministic predictions. Such a model is also evaluated using the RMSE metrics. Probabilistic metrics (e.g., log-score, CRPS; see Baggio and Muzy (2024)) will be explored in a future study.
In section 4-4.2.2 we study the performance of the model in forecasting extremes. The occurrence or non-occurrence of an extreme wind value is given by a binary variable $I^{s}_{d,t}$ such that $I^{s}_{d,t}=1$ if the event occurs, while $I^{s}_{d,t}=0$ otherwise. Then the quality of binary forecasts ${\widehat{I}}^{s}_{d,t}$ needs to be evaluated by a relevant metric. This is usually referred to as a “classification problem” (as opposed to a “regression problem” as formerly considered). In that case, one can use a variety of metrics (or skill scores), depending upon which aspect of the prediction one wants to optimize. We can refer the reader to Jolliffe (2004) (see also Wilks (2019)) for a comprehensive review on the topic. Among all possible scores, the Peirce Skill Score (PSS) is a widely used metric for evaluating the performance of a binary classification model, particularly in the context of severe weather prediction. It is defined as:
[…]
where HR stands for “Hit Rate” and FA for “False Alarm Rate”. These rates are derived from the contingency table of predictions, where the $N$ forecast/observation pairs $({\widehat{I}}^{s}_{d,t},I^{s}_{d,t})$ are divided in true positives TP, false positives FP, false negatives FN and true negatives TN (we thus have TP+FP+FN+TN=N). More specifically, $\text{HR}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ corresponds to the fraction of correctly predicted positive events, while $\text{FA}=\frac{\text{FP}}{\text{TN}+\text{FP}}$ is the fraction of wrongly predicted, i.e. predicted as positive, negative events. Unlike accuracy-based scores, PSS accounts for both missed events (false negatives) and false alarms (false positives), providing a balanced assessment of quantifying model performance. Moreover, this score is particularly suitable for forecasting extremes as it weights more strongly correct yes forecasts of rare events, thus discouraging artificial distortions towards the more common class. The PSS ranges from $-1$ to $1$, with $\text{PSS}=1$ corresponding to a perfect forecast and $\text{PSS}=0$ corresponding to unskilled, i.e., purely random, forecast. This means that a negative PSS value reflects prediction performance that is worse-than-random. Another score commonly used when forecasting rare events (Jolliffe, 2004) is the Critical Success Index (CSI), also known as Threat Score. It is defined as:
$\text{CSI}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}\,,$ (13)
corresponding to the ratio between the times $I^{s}_{d,t}=1$ has been forecasted correctly and the total number of times the event either happened, has been forecasted or both. The maximum value for CSI is 1, while the lowest possible value is 0. As TN are not considered in 13, this metric is not artificially inflated by correct forecasts of the more common ”no” event. However, contrary to PSS, it is not an equitable metric, which makes comparisons between binary events with different frequency of occurrence not straightforward. In the following, we refer mostly to the PSS when assessing performance, however, we will show some results in terms of CSI as well.
3.6 Baseline models
In order to compare the performance of previous neural network models with alternative approaches, we consider various simple or natural baselines. The first model we consider corresponds to the raw output directly provided by the NWP regional model, namely the AROME model. For a given site $s$, some given day $d$ and time $t$, at the lead time $h$, we use the wind speed values from the closest grid point in AROME at time $t+h$. This will be referred to as the “Raw AROME” prediction. Although, as discussed in the introduction, such raw predictions can often exhibit some bias and can be improved through different post-processing approaches, it is the simplest baseline model one can use for reference. Along the same line we also consider “Raw ARPEGE” as the prediction provided by the global ARPEGE model at the grid point closest to the considered site. Given that AROME predictions are downscaled from ARPEGE, it is expected that AROME will deliver superior performance over ARPEGE. Another model that is usually considered as a reference model for surface wind speed is the so-called “Persistence”, which consists in considering the last measured value as the best forecast. Mathematically, this amounts to assuming that the wind speed is a Martingale random process. Finally, for the sake of completeness, we consider the class of linear models as last comparative models. Such models can notably be considered as the most basic neural network with a single fully connected layer with a linear output and which takes the past observed wind speed values as input. This model will be referred to as ${\cal L}$ (for linear) model.
@@ 4.1.1 Comparison with basic models: raw AROME prediction and persistence models
Table 3: RMSE relative improvement of ${\mathcal{M}}$ as respect to baseline models at horizons $1$ and $6$ hours
@@
Figure 3: Comparison of ${\mathcal{M}}$ model performance as respect to various baseline models described in the text. The RMSE over the test sample is reported for each prediction horizon.
The comparative performances in terms of RMSE of the ${\mathcal{M}}$ model and the various baseline models are reported in Fig. 3. These results correspond to the RMSE evaluated over the test set for all 278 stations in the MeteoNet dataset for prediction horizons $h=[1,2,3,4,5,6]$ hours.One can see that the ${\mathcal{M}}$ model significantly outperforms all baselines for all horizons, as highlighted in Table 3 for horizons $1$ and $6$ hours. We see that Linear ($\cal L$) and Persistence models based on station data provide good performances at very short time horizons but quickly become less reliable as the forecasting horizon increases. On the other hand, as anticipated in the discussion of Sec. 3-3.1, we observe that raw NWP predictions (AROME and ARPEGE) errors hardly vary with horizon. We also observe that, as expected, the regional, high-resolution AROME model performs better than the global ARPEGE model. AROME becomes better than a regression of station data at horizons greater that 3 $h$. However, both models are surpassed by ${\mathcal{M}}$, even at largest time horizons. This clearly demonstrates how ${\mathcal{M}}$, which can be considered as a kind of post-processing method, can greatly improve the prediction accuracy of raw NWP forecasts by leveraging multiple sources of input data and correcting bias inherent to these models.
Fig. 4 illustrates model predictions for the specific case of two ground station sites in Corsica: Ajaccio and Bonifacio. In Ajaccio, we observe a pronounced annual oscillation related to the coastal breeze regime, whereas in Bonifacio, wind speeds are more intermittent and wind gusts more intense. In both subplots, we present the values of 6-hour ahead forecasts from the AROME and ${\mathcal{M}}$ models at a fixed time of day $t_{d}=$ 1100 UTC (the corresponding initial time is therefore $t=t_{d}-h$ = 0500 UTC) for all days over the 2016–2018 period. For the sake of clarity, the plots include all three sub-periods (training, validation, and test subsets), which may result in slightly overestimated model performance compared to its actual values. We see that both model predictions closely align with ground truth values at both locations for all time periods. It can be however remarked that AROME model slightly underestimates wind speeds, particularly during winter, at both sites (in Fig. 4, orange squares are more likely to be observed under black solid lines that blue circles). These observations highlight the need for some bias correction in raw numerical weather prediction (NWP) outputs, reinforcing the importance of post-processing techniques in improving forecast accuracy.
It is important to note that, beyond the average results presented in Fig. 3, the RMSE-based forecasting performance of the ${\mathcal{M}}$ model varies significantly across stations. This variability stems from the strong spatial variability in site-specific weather conditions, particularly mean wind speed. As shown in Fig. 5, there is a clear linear dependence between RMSE and mean wind speed estimated at each site over the full 3 years period, with higher mean value leading to larger prediction error.
Figure 4: Illustration of the ${\mathcal{M}}$ model and the raw AROME forecasting performances in Ajaccio and Bonifacio for $h=6$ h ahead prediction of wind speed value at time of the day $t_{d}=t+h=1100UTC$. At first sight performances are comparable but one can see that the regional NWP model is slightly biased towards small wind speed values. One can also observe strong seasonal effects in Ajaccio where breeze regimes are much more important than in Bonifacio site.
@@
Figure 5: RMSE associated with the ${\mathcal{M}}$ prediction as a function of the station mean wind speed value (estimated over 3 years) $\overline{V}^{s}$ for all sites for horizons $h=1$ hour and $h=6$ hours.
@@ 4.1.2 Comparison with single branch models
In Fig. 6 are displayed, as a function of the horizon, the RMSE performances of the ${\mathcal{M}}$ model comparatively to ${\mathcal{M}}_{AR}$, ${\mathcal{M}}_{AP}$ and ${\mathcal{M}}_{GS}$ models that respectively process AROME, ARPEGE and Ground Station input data without other branches. We first see that ${\mathcal{M}}$ predictions are the best among all alternatives for all times horizons. Let us also observe, by comparing the results of Figs 3 and 6, that post-processing NWP inputs, whether AROME or ARPEGE, with a neural network model significantly reduces their error. Specifically, the RMSE for ${\mathcal{M}}_{AR}$ predictions is reduced by approximately 27 % compared to raw AROME predictions, while the RMSE for ${\mathcal{M}}_{AP}$ predictions is reduced by about 36 % compared to raw ARPEGE predictions. ${\mathcal{M}}_{AR}$ model associated with AROME remains better than ${\mathcal{M}}_{AP}$, the one associated with ARPEGE, even if the difference between the two models is reduced as compared to the raw predictions. In both models, as before and as expected from discussion in Sec. 3-3.1, the error only slightly depends on the time horizon. On the other hand, as remarked previously, the performance of ${\mathcal{M}}_{GS}$, the model relying exclusively on past observations at ground station and its surrounding sites, strongly depends on the horizon and clearly decreases as the horizon increases.
Figure 6: Comparison of ${\mathcal{M}}$ model performance as respect to ${\mathcal{M}}_{AR}$, ${\mathcal{M}}_{AP}$ and ${\mathcal{M}}_{GS}$ models. The RMSE over the test sample is reported for each prediction horizon.
@@ 4.1.3 Feature importance analysis
In the previous paragraph, we have seen that the model ${\mathcal{M}}$, which integrates all three data sources, consistently outperforms the best individual model at every forecast horizon. The model leverages the strengths of each model at each time horizon, optimizing performance across all forecasts. In order to assess the contribution of each feature to the reduction in final RMSE, we adopt the approach outlined in Rasp and Lerch (2018), which is based on a method originally proposed by Breiman (2001) for random forest models. This method simply consists for a given input feature or group of features, in randomly shuffling the corresponding data among a given mini-batch sample. The main advantage of this method is its relative simplicity and the fact that it does not require retraining the model. This is not the case of the approach used by Marcille et al. (2024), which involves analyzing the impact of a specific subset of features by removing it from the original input and run a new training phase within this configuration.
@@
In Fig. 7 are reported the importance score at horizons $h=1$, $h=3$ and $h=6$ hours for each feature in the ${\mathcal{M}}$ model computed on the test sample for the whole dataset. One can see that NWP and station wind components inputs are, by far, the most important features. Among other features, the relative humidity predicted by AROME model and the temporal features (date, hour of the day) have the most significant impact of the prediction performances. As far as horizons are concerned, there is no clear trend in AROME feature importance whereas ARPEGE wind feature become more important as the horizon increases. This is not surprising since ARPEGE is a global model that is designed to capture large scale features several days ahead. On the other hand, past station features become less important for longer horizons as compared to shorter ones. Models based on observed past weather conditions at the target site and its surrounding ground stations are more likely to be useful at small horizons than for longer ones. It is noteworthy that for both AROME and ARPEGE temperature data have more impact on the prediction quality than pressure data. This may be reminiscent of the correlation of the wind components with thermal breezes as observed by Marcille et al. (2024) for the meridional wind at the coastal station they analyzed.
@@ 4.1.4 The benefits of a probabilistic approach
Figure 8: Comparison of ${\mathcal{M}}$ ($\blacksquare$) and ${\mathcal{M}}^{\prime}$ model ($\bullet$) RMSE for all stations as a function of the forecasting horizon.
In Fig. 8, we compare the RMSE observed on the test set for the ${\mathcal{M}}$ model (blue ($\blacksquare$) symbols) and the ${\mathcal{M}}^{\prime}$ model (red ($\bullet$) symbols). Recall that the latter is designed to predict the parameters of the M-Rice probability distribution for the wind speed over the next six hours. At each time step, one can then compute, for any given forecasting horizon, the conditional mean using Eq. (20) and then the resulting MSE over the validation or the test samples. As it can be observed in Fig. 8, whatever the considered horizon the probabilistic model (${\mathcal{M}}^{\prime}$) provides slightly better forecasts than the deterministic version (${\mathcal{M}}$). A priori, the model designed to optimize directly the empirical MSE (${\mathcal{M}}$) is expected to provide better results in terms of MSE as compared to any other one.
It appears, from Fig. 8, that parametrizing directly the conditional mean is less efficient that trying to estimate it through the probability law it is associated with. Let us observe that this is even more true as the forecast horizon increases. Moreover, as advocated previously, probabilistic forecasts offer valuable insights beyond a single-point prediction. They provide the entire probability distribution, offering the possibility to compute various risk measures and facilitating decision-making tailored to specific objectives. Even for single-point prediction, as considered in this paper, if one uses different metrics, the probability law can provide more efficient prediction than the MSE. For example, if one aims at minimizing another metrics instead of the MSE (e.g. the mean absolute error), there is no need to perform another training of the model since it suffices to compute, for each station and for each time $t$, the value the optimizes this metrics as the best forecast instead of the conditional mean value (e.g., the conditional median). In Sec. 4-4.2-4.2.2, we will see that in the case of extreme wind occurrence prediction, high quantile values threshold exceedances provide a very good predictor which outperforms significantly the deterministic predictions.

### v2

@@ 4.3 Baseline models
All models presented in this study, including both deterministic and probabilistic architectures, were implemented and trained using the Keras API with TensorFlow (Abadi et al., 2015; Chollet et al., 2015). The TensorFlow framework provides efficient support for GPU acceleration and large-scale data pipelines, which was essential for training the models on the MeteoNet dataset and for rapid evaluation during fine-tuning and validation phases. All training were performed using a single Nvidia V100 GPU. The total training time for models $\mathcal{M}$ and $\mathcal{M^{\prime}}$, that is, the most computationally demanding among models considered in this study, is approximately 3 hours (about 220 seconds per epoch with $\approx 50$ epochs needed for convergence). As typical for NNs, training is performed only once. Afterward, the model can generate predictions on a new batch of data in just a few seconds ( $\approx\,5\,s$ including the time needed for model loading and set-up).
@@ 5.1.1 Comparison with basic models: raw AROME prediction and persistence models
Table 3: RMSE relative improvement of $\mathcal{M}$} with respect to baseline models at horizons $1$ and $6$ hours.
@@
Figure 3: Comparison of the performance of the $\mathcal{M}$} model (symbols ($\blacksquare$)) against various baseline approaches described in Secs. 4.2 and 4.3. The Root Mean Square Error (RMSE), expressed in $ms^{-1}$ and defined in Eq. (5), is computed over the test dataset and reported for each forecast horizon ranging from 1 to 6 hours.
[…]
Figure 4: Illustration of the $\mathcal{M}$} (symbols ($\bullet$)) model and the raw AROME forecasting performances ($\blacksquare$) in Ajaccio and Bonifacio for horizon $h$ == 66 hours ahead prediction of wind speed value at valid time $_{h} == tt$ ++ $h$ == $1100$ UTC each day. Measured wind speed value as indicated by the black solid lines. Panels (a) and (c) display daily forecasts over the full 2016–2018 period, while panels (b) and (d) provide a zoomed view of a representative 4-month sub-period (15 January 2016 to 15 May 2016) to enhance visual clarity. Although the overall performance of the two models appears comparable in (a) and (c), the raw AROME forecasts exhibit a slight bias toward lower wind speeds. One can also observe strong seasonal effects in Ajaccio where breeze regimes are much more important than in Bonifacio. The zoomed views in (b) and (d) confirm that the $\mathcal{M}$} model outperforms the raw AROME forecasts quite frequently, on a day-to-day basis, indicating consistent statistical superiority.
[…]
5.1.2 Model performance as a function of site mean wind speed and across valid times
This section investigates how the forecasting accuracy of model $\mathcal{M}$} varies with local wind conditions and time of day, highlighting the influence of mean wind speed and diurnal cycles on prediction errors.
@@ 5.1.2 Model performance as a function of site mean wind speed and across valid times
Figure 5: RMSE (in $ms^{-1}$) associated with the $\mathcal{M}$} prediction as a function of the station mean wind speed value (estimated over 3 years) $\overline{V}^{s}$ for all sites for horizons $h$ = 11 hour (symbols ( $\bullet$)) and $h$ = 66 hours (symbols ($\bigstar$)).
It is first important to note that, beyond the average results presented in Fig. 3, the RMSE-based forecasting performance depends of the $\mathcal{M}$} model varies significantly across stations. This variability stems from the strong spatial variability in site-specific weather conditions, particularly mean wind speed. As shown in Fig. 5, there is a clear linear dependence between RMSE and mean wind speed estimated at each site over the full 3 years period, with higher mean value leading to larger prediction error.
Figure 6: Diurnal variation of (a) RMSE and (b) relative RMSE (RRMSE) as a function of the valid time for model $\mathcal{M}$} predictions at horizons $h$ ranging from 1 hour to 6 hours. Results are averaged over the 278 stations of the MeteoNet dataset. The RMSE (panel a) slowly increases with valid time whereas the RRMSE (panel b), normalized by the mean wind speed, shows a U-shaped pattern.
Another key factor that likely influences the model’s forecasting performance is the valid time—the targeted time of day for which the prediction is issued, as defined in Sec. 3. Figure 6 illustrates how prediction errors, expressed as RMSE and Relative RMSE, vary with valid time. Here, valid time is the sum of the initial time and the forecast horizon, with results shown for horizons from $h=1$ h to $h=6$ h and averaged over the 278 ground stations in the dataset. In panel (a), RMSE generally increases with forecast horizon: for a given valid time, earlier-issued forecasts tend to be less accurate. RMSE also shows a slight increase with valid time up to around 16:00–17:00 UTC, after which it gradually decreases. This peak may reflect the fact that wind regimes typically become more turbulent and intermittent from early morning to mid-afternoon, reducing predictability. Nonetheless Panel (b) presents Relative RMSE (RRMSE), computed by normalizing the RMSE at each valid time by the corresponding mean wind speed. This normalization highlights a pronounced U-shaped pattern: forecast skill as measured relatively to the mean velocity is lowest during the night and early morning, peaks around 14:00 UTC and then declines again in the late afternoon. This behavior primarily reflects the regional diurnal wind cycle, which follows a bell-shaped profile: winds are generally weak at night and in the early morning, then intensify through the afternoon under the influence of thermally driven sea-breeze circulations (Dai and Deser, 1999), sometimes reinforced by synoptic flows such as the Mistral or Tramontane in the French Mediterranean.
@@ 5.1.3 Comparison with single branch models
In Fig. 7 are displayed, as a function of the horizon, the RMSE performances of the $\mathcal{M}$} model comparatively to ${\mathcal{M}}_{AR}$, ${\mathcal{M}}_{AP}$ and ${\mathcal{M}}_{GS}$ models that respectively process AROME, ARPEGE and Ground Station input data without other branches. We first see that $\mathcal{M}$} predictions are the best among all alternatives for all times horizons. Comparing the results of Figs. 3 and 7 shows that post-processing NWP inputs, whether AROME or ARPEGE, with a neural network model significantly reduces their error. Specifically, the RMSE for ${\mathcal{M}}_{AR}$ predictions is reduced by approximately 27 % compared to raw AROME predictions, while the RMSE for ${\mathcal{M}}_{AP}$ predictions is reduced by about 36 % compared to raw ARPEGE predictions. The ${\mathcal{M}}_{AR}$ model associated with AROME remains better than ${\mathcal{M}}_{AP}$ associated with ARPEGE, even if the difference between the two models is reduced as compared to the raw predictions. In both models, as before and as expected from the discussion in Sec. 3, the error only slightly depends on the time horizon. On the other hand, as remarked previously, the performance of ${\mathcal{M}}_{GS}$, the model relying exclusively on past observations at ground station and its surrounding sites, strongly depends on the horizon and clearly decreases as the horizon increases.
Figure 7: Comparison of the performance of the $\mathcal{M}$} model (symbols ($\blacksquare$)) as respect to the ${\mathcal{M}}_{AR}$ (symbols ()), ${\mathcal{M}}_{AP}$ (symbols ()) and ${\mathcal{M}}_{GS}$ models (symbols ($\bigstar$)). The RMSE expressed in $ms^{-1}$ is computed over the test dataset and reported for each forecast horizon ranging from 1 to 6 hours.
@@ 5.1.4 Feature importance analysis
In the previous paragraph, we have seen that the model $\mathcal{M}$}, which integrates all three data sources, consistently outperforms the best individual model at every forecast horizon. The model leverages the strengths of each model at each time horizon, optimizing performance across all forecasts. In order to assess the contribution of each feature to the reduction in final RMSE, we adopt the approach outlined in Rasp and Lerch (2018), which is based on a method originally proposed by Breiman (2001) for random forest models. This method simply consists, for a given input feature or group of features, in randomly shuffling the corresponding data among a given mini-batch sample. The main advantage of this method is its relative simplicity and the fact that it does not require retraining the model. This is not the case of the approach used by Marcille et al. (2024), which involves analyzing the impact of a specific subset of features by removing it from the original input and run a new training phase within this configuration.
@@
In Fig. 8 are reported the importance score $S_{F}$ at horizons $h=1$, $h=3$ and $h=6$ hours for each feature in the $\mathcal{M}$} model computed on the test sample for the whole dataset. One can see that NWP and station wind components inputs are, by far, the most important features. Among other features, the relative humidity predicted by the AROME model and the temporal features (date, hour of the day) have the most significant impact of the prediction performances. As far as horizons are concerned, there is no clear trend in AROME feature importance whereas ARPEGE wind feature become more important as the horizon increases. This is not surprising since ARPEGE is a global model that is designed to capture large scale features several days ahead. On the other hand, past station features become less important for longer horizons as compared to shorter ones. Models based on observed past weather conditions at the target site and its surrounding ground stations are more likely to be useful at small horizons than for longer ones, as highlighted by Fig. 7. It is noteworthy that for both AROME and ARPEGE temperature data have more impact on the prediction quality than pressure data. This may be reminiscent of the correlation of the wind components with thermal breezes as observed by Marcille et al. (2024) for the meridional wind at the coastal station they analyzed.
Beyond its methodological interest, the feature importance analysis provides valuable insights for stakeholders and system operators. By quantifying the impact of each input feature on forecast accuracy, the importance scores help identify which data sources are most critical for reliable predictions. For example, the dominant role of NWP wind components and station observations confirms the operational relevance of maintaining high-quality measurements and forecasts from these sources. Similarly, the increasing importance of ARPEGE features at longer horizons suggests that global-scale forecasts are particularly useful for planning decisions further ahead in time. This information can guide decisions on sensor deployment, data validation, and resource allocation. It also supports transparency and trust in the model by offering a clear, quantifiable link between input data and output performance. Moreover, such analysis can be leveraged to design more compact and efficient models by identifying and potentially excluding less informative inputs. This can reduce the number of model parameters, simplify training, and improve generalization. Simpler models are also more interpretable and computationally efficient, making them better suited for real-time forecasting applications and easier to maintain.
@@ 5.1.5 The benefits of a probabilistic approach
Figure 9: Comparison of model $\mathcal{M}$} (symbols ($\blacksquare$)) and model $\mathcal{M}$$^{\prime}$ (symbols ($\bullet$)) performance in their ability to provide a single point prediction. RMSE, in $ms^{-1}$, is plotted as a function of the forecasting horizon in hours.
In Fig. 9, we compare the RMSE observed on the test set for the $\mathcal{M}$} model (blue ($\blacksquare$) symbols) and the $\mathcal{M}$$^{\prime}$ model (red ($\bullet$) symbols). Recall that the latter is designed to predict the parameters of the M-Rice probability distribution for the wind speed over the next six hours. At each time step, one can then compute, for any given forecasting horizon, the conditional mean using Eq. (20) and then the resulting MSE over the validation or the test samples. As it can be observed in Fig. 9, whatever the considered horizon the probabilistic model ($\mathcal{M}$$^{\prime}$) provides slightly better forecasts than the deterministic version ($\mathcal{M}$}). To ensure the statistical significance of these results, we applied the Wilcoxon signed-rank test (Wilcoxon, 1992), a non-parametric test for paired data that evaluates whether the median difference between two related samples, in this case, the forecast errors from $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$, is significantly different from zero. The test confirms that $\mathcal{M}$$^{\prime}$ is more accurate, as the obtained p-values are very close to zero ($p<10^{-20}$ for all horizons).
A priori, the model designed to optimize directly the empirical MSE ($\mathcal{M}$}) is expected to provide better results in terms of MSE as compared to any other one. Fig. 9 indicates that parametrizing directly the conditional mean is less efficient that trying to estimate it through the probability law it is associated with. This is even more true as the forecast horizon increases. Moreover, as advocated previously, probabilistic forecasts offer valuable insights beyond a single-point prediction. They provide the entire probability distribution, offering the possibility to compute various risk measures and facilitating decision-making tailored to specific objectives. Even for single-point prediction, as considered in this paper, if one uses different metrics, the probability law can provide more efficient prediction than the MSE. For example, if one aims at minimizing another score instead of the MSE (e.g. the mean absolute error), there is no need to perform another training of the model since it suffices to compute, for each station and for each time $t$, the value that optimizes this score as the best forecast instead of the conditional mean value (e.g., the conditional median). Section 5.2.3 will show that in the case of extreme wind occurrence prediction, high quantile values threshold exceedances provide a very good predictor which outperforms significantly the deterministic predictions.